OpenStudy (anonymous):

Solve. √x+ 7 + √x =2

OpenStudy (anonymous):

\[\sqrt{x+7} + \sqrt{x} = 2\]

OpenStudy (anonymous):

First thought is to use logs, do you know how to use those?

OpenStudy (anonymous):

Hrm.... well you could "square" both sides of the equation.\[\huge (\sqrt{x+7}+\sqrt x)^2 = 2^2\]

OpenStudy (anonymous):

Would it be no solution?

OpenStudy (anonymous):

I dunno, i haven't done the math yet. did you do the math that i posted?